OpenStudy (anonymous):

Tara's car accelerated from 10 miles per hour to 60 miles per hour in 12 seconds. What is the approximate rate of change in the miles per hour that Tara's car accelerated?

OpenStudy (athecool):

Ask @mathmale to help he's awesome!

OpenStudy (athecool):

@mathmale

OpenStudy (anonymous):

(end velocity - start velocity)/time elapsed

OpenStudy (anonymous):

a=(v-u)/t u=initial velocity v=final velocity t=time a=acceleration convert to meters/second 60=26.8224 10=4.4704 a(rate of change in meters/second^2) = (26.8224-4.4704)/12 =1.86266666667 ms^(-2)

OpenStudy (anonymous):

or 1500 miles/h^2

OpenStudy (anonymous):

If the question has been answered, please close it, @maryweisbeck